Lesson : Periodic Motion



1. A device that is used for medical diagnosis oscillates at a frequency of 6.7 X 10⁶ Hz. How much time does each oscillation take? What is also its angular frequency?



2. Determine the displacement of an object that is undergoing SHM whose angular


frequency is 20°/ sec and its amplitude is 300mm at t = 1.5 sec.



3. Write down at least three (3) real life situations where periodic motion is applicable.

1
Expert's answer
2021-11-23T10:31:14-0500

1. "\nu=1/T\to T=1/\nu=1/(6.7\cdot10^6)=1.49\cdot10^{-7}\ (s)"

"\omega =2\pi\nu=2\cdot3.14\cdot6.7\cdot10^6=42076000\ (rad/s)"


2. "20 \ °/s=0.35\ rad/s"


For SHM "s=A\sin\omega t" . If "t=0" then "s=0" .


"0.3=A\sin0.35 t\to0.3=A\sin0.35\cdot1.5\to A=0.6\ (m)"


"s=0.6\sin0.35t"


"\Delta s=0.3\ (m)"


3. The rocking chair, the tuning fork, the swing, the pendulum in the clocks, car shock absorber, musical instruments, hearing, cradle, bungee jumping.